Were the expected and observed results similar?

4. Were the expected and observed results similar? Do your results match the prediction for a population in Hardy-Weinberg equilibrium? If not suggest an explanation. 

5. What do you expect to happen to the genotype and allele frequencies of the population after running this simulation for 20 generations? 
Answer #1

Qn4.

RR = 7, Rr=18, rr=15

So the population size is 40.

Frequency of R allele = [7+(18/2)] /40 = 0.4

Frequency of r allele = [15+(18/2)]/40 =0.6

We know according to Hardy-Weinberg equilibrium, p+q=1

where dominant allele Frequency= p , recessive allele Frequency = a

(p+q)2=1

p2+2pq+q2=1

p2= Homozygous dominant genotype freq.

2pq= Heterozygous genotype frequency

q2= Homozygous recessive genotype Frequency

observed Expected 0-E 0-E) (0-E)% 7 7 o RR O 18 19 Ro - 1 005 15 14 - rr o.of 40 40 = 0.12 6-E)? E Degrees of freedom 3-1 22

Qn5.

After 20 generation the allele Frequencies and genotype will remain same. Because the population is in Hardy-Weinberg equilibrium.

According to Hardy-Weinberg principle, a population which is in Hardy-Weinberg equilibrium, the allele Frequency remains unchanged generation after generation.